how do i solve 4y=1?
Answer by jim_thompson5910(35256)

4y=1 ... Start with the given equation


4y/4=1/4 ... Divide both sides by 4 to isolate 'y'


y = 1/4


So the solution is y = 1/4


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim